OpenStudy (anonymous):

Write an equation of a line that passes through (9,-5) and (3,-5).

OpenStudy (cp9454):

y+5= 0

OpenStudy (goldphenoix):

Does it say in slope-intercept form? (y = mx + b)

OpenStudy (anonymous):

No, (y2-y1)/(x2-x1)

OpenStudy (anonymous):

use this.. (y-y1)/(x-x1)=(y2-y1)/(x2-x1)

OpenStudy (goldphenoix):

That's for finding the slope.